OpenStudy (loser66):

Please, help \[\int\dfrac{1}{3sinx+1}dx\]

OpenStudy (mathmale):

Faiq: Students generally must learn how to solve problems themselves. You might suggest their use of www.integral-calculator.com only after your student has made and shared some effort to integrate on his own.

OpenStudy (mathmale):

Thoughts: Have you tried integration by parts? Any chance that you could simplify the integration by multiplying numerator and denominator of \[\frac{ 1 }{ 3\sin x + 1 }\]

imqwerty (imqwerty):

hint: multiply and divide by 3sinx-1

OpenStudy (mathmale):

by (3 xin x - 1)?

OpenStudy (mathmale):

Great minds think alike.

imqwerty (imqwerty):

^yes :)

OpenStudy (mathmale):

:)

OpenStudy (faiqraees):

Oh maybe we can do this \[\large \rm \int \frac{1}{3sinx+1} *1\]

OpenStudy (mathmale):

\[\large \rm \int\limits \frac{1}{3sinx+1} *1\] is incomplete (and not proper) unless you include that differential, "dx."

OpenStudy (mathmale):

\[\large \rm \int\limits \frac{1}{3sinx+1} *1dx\]

OpenStudy (faiqraees):

but wouldn't we get a much more complex integral in its solution by parts?

imqwerty (imqwerty):

by doing this \(\large \rm \int\limits \frac{1}{3sinx+1} *1dx\) we'll encounter this-\[\int\limits_{}^{}\frac{ 3cosx }{ (3sinx+1)^2 }dx\] which will get complex

imqwerty (imqwerty):

it must be this*\(\int\limits_{}^{}\frac{x (3cosx )}{ (3sinx+1)^2 }dx\)

imqwerty (imqwerty):

no that is alright^^ but when you solve it further using integration by parts then you'll encounter this^

OpenStudy (loser66):

Unfortunately, we go nowhere there.

OpenStudy (loser66):

(3sinx +1) (3sinx -1) = 9sin^2 x -1

OpenStudy (loser66):

then??

OpenStudy (loser66):

If we multiple by 3sin x +1, the integrand becomes \(\dfrac{3sinx +1}{9sin^2x +6sin x+1}\) , then??

OpenStudy (loser66):

No outcome!!

OpenStudy (loser66):